# RLHF and IIA: Perverse Incentives

### 02 Dec 2023 v1

@@ A.2 DPO
$P_{\pi_{r}}(x)=\frac{P_{\overline{\pi}}(x)e^{r(x)N/\beta}}{\sum_{x^{\prime}\in\mathcal{M}}P_{\overline{\pi}}(x^{\prime})e^{r(x^{\prime})N/\beta}}.$ (28)
@@
$-\sum_{(\mathcal{Y},y)\in\mathcal{D}}\left[\ln\left(\frac{e^{r(y)}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{r(y^{\prime})}}\right)\right]=-\sum_{(\mathcal{Y},y)\in\mathcal{D}}\left[\ln\frac{\left(P_{\pi_{r}}(y)/P_{\overline{\pi}}(y)\right)^{\beta/N}}{\sum_{y^{\prime}\in\mathcal{Y}}\left(P_{\pi_{r}}(y^{\prime})/P_{\overline{\pi}}(y^{\prime})\right)^{\beta/N}}\right].$ (29)
Thus the value $\mathcal{L}_{\mathrm{reward}}(r)$ equals the value $\mathcal{L}_{\mathrm{DPO}}(\pi_{r})$. Note also that $r\mapsto\pi_{r}$ is surjective.
Suppose $\widehat{r}$ is the minimizer for $\mathcal{L}_{\mathrm{reward}}$. Then $\pi_{\widehat{r}}$ is the optimal policy obtained by minimizing $\mathcal{L}_{\mathrm{policy}}(\cdot|\widehat{r})$. If $\pi_{\widehat{r}}$ is not optimal for $\mathcal{L}_{\mathrm{DPO}}$, then there exists another policy $\pi^{\prime}$ that obtains a strictly lower loss $\mathcal{L}_{\mathrm{DPO}}(\pi^{\prime})$. Since $r\mapsto\pi_{r}$ is surjective, there exists a reward function $r^{\prime}$ such that $\pi^{\prime}=\pi_{r^{\prime}}$. For example, we can take $r^{\prime}(x)=\frac{\beta}{N}\ln\frac{P_{\pi^{\prime}}(x)}{P_{\overline{\pi}}(x)}$, which achieves a lower reward loss $\mathcal{L}_{\mathrm{reward}}(r^{\prime})=\mathcal{L}_{\mathrm{DPO}}(\pi_{r^{\prime}})<\mathcal{L}_{\mathrm{DPO}}(\pi_{\widehat{r}})=\mathcal{L}_{\mathrm{reward}}(\widehat{r})$, a contradiction.
Similarly, if $\pi^{*}$ is optimal for $\mathcal{L}_{\mathrm{DPO}}$, the corresponding reward $r^{*}(x)=\frac{\beta}{N}\ln\frac{P_{\pi^{*}}(x)}{P_{\overline{\pi}}(x)}$ must be optimal for $\mathcal{L}_{\mathrm{reward}}$. The corresponding optimal policy $\pi_{r^{*}}$ for $\mathcal{L}_{\mathrm{policy}}(\cdot|r^{*})$ then satisfies
$P_{\pi_{r^{*}}}(x)\propto P_{\overline{\pi}}(x)e^{\frac{\beta}{N}\ln\frac{P_{\pi^{*}}(x)}{P_{\overline{\pi}}(x)}N/\beta}=P_{\pi^{*}}(x),$
as desired.
@@ A.3 Inclusive Learning
Similar to the proof for Lemma 9, we define
[…]
as the set of indices of data in $\mathcal{D}=\{(\mathcal{Y}^{(n)},y^{(n)})\}_{n=1}^{N}$ that contain both categories of messages. Note that an alternative definition of $\mathcal{I}$ is
=$\mathcal{I}=\{n\in[N]\,|\,\mathcal{Y}^{(n)}\cap\mathcal{M}_{1}\neq\emptyset,\mathcal{Y}^{(n)}\cap\mathcal{M}_{2}\neq\emptyset\}.$
Then we have the following lemma:
@@
Under Assumptions 1 and 3, if $K\geq 2$, $\widehat{\theta}$ minimizes $\mathcal{L}_{\mathrm{inclusive}}(\widehat{\pi}_{\theta})$, for all fixed $\mathcal{M}_{1}$ and $\mathcal{D}$ that satisfy
  1. 1.
    $\exists n,n^{\prime}\in\mathcal{I}$ such that $y_{n}\in\mathcal{M}_{1}$ and $y_{n^{\prime}}\in\mathcal{M}_{2}$,
  2. 2.
    |$|\{n\in\mathcal{I}\,|\,y^{(n)}\in\mathcal{M}_{2}\}|>\max\{1+\beta,2\beta\ln\frac{p_{*}(1)}{p_{*}(2)}\}$,
then
@@
First, we notice that $\mathcal{L}_{\mathrm{inclusive}}(\widehat{\pi}_{\theta})$ is strongly convex for $\beta>0$ and $\mathcal{L}_{\mathrm{inclusive}}(\widehat{\pi}_{\theta})\to\infty$ as $\|\theta\|_{2}\to\infty$, thus a minimizer for $\mathcal{L}_{\mathrm{inclusive}}(\widehat{\pi}_{\theta})$ exists and is unique. As a shorthand, for a policy $\pi$, we use $\mathbf{p}_{\pi}=(p_{\pi}(1),p_{\pi}(2))$ to denote the two-dimensional vector on the probability simplex $(P_{\pi}(\mathcal{M}_{1}),P_{\pi}(\mathcal{M}_{2}))$, and $m_{z}$ to denote $|\mathcal{M}_{z}|$. Letting $m=m_{2}/m_{1}$, the inclusive loss can be written in terms of $\mathbf{p}_{\pi}$ and $\mathbf{p}_{\overline{\pi}}$ as
$\mathcal{L}_{\rm IL}(\mathbf{p}_{\pi})=\tilde{\mathcal{L}}_{N,m}(\mathbf{p}_{\pi})+\beta\cdot\mathbf{d}_{\mathrm{KL}}(\mathbf{p}_{\pi}\|\mathbf{p}_{\bar{\pi}}),$ (30)
where ~$\tilde{\mathcal{L}}_{N,m}(\mathbf{p}_{\pi})$ is the unregularized inclusive learning loss
~$\tilde{\mathcal{L}}_{N,m}(\mathbf{p}_{\pi})$ $=$ $-\sum_{n=1}^{N}\ln\frac{p_{\pi}(1)/m_{1}\cdot\mathbbm{1}(y_{n}\in\mathcal{M}_{1})+p_{\pi}(2)/m_{2}\cdot\mathbbm{1}(y_{n}\in\mathcal{M}_{2})}{k_{1}^{(n)}p_{\pi}(1)/m_{1}+k_{2}^{(n)}p_{\pi}(2)/m_{2}}$ (31)
@@
$=$ $-\sum_{n=1}^{N}\ln\frac{mp_{\pi}(1)\cdot\mathbbm{1}(y_{n}\in\mathcal{M}_{1})+p_{\pi}(2)\cdot\mathbbm{1}(y_{n}\in\mathcal{M}_{2})}{mk_{1}^{(n)}p_{\pi}(1)+k_{2}^{(n)}p_{\pi}(2)}.$
Substituting $p_{\pi}(1)$ with $1-p_{\pi}(2)$, the above loss becomes a function of $p_{\pi}(2)$. To simplify the exposition, we use $\zeta$ to denote $p_{\pi}(2)$:
~$\tilde{\mathcal{L}}_{N,m}(\mathbf{p}_{\pi})=f_{N,m}(p_{\pi}(2))\qquad\mathbf{d}_{\mathrm{KL}}(\mathbf{p}_{\pi}\|\mathbf{p}_{\overline{\pi}})=(1-\zeta)\ln\frac{1-\zeta}{p_{*}(1)}+\zeta\ln\frac{\zeta}{p_{*}(2)}$
with
$f_{N,m}(\zeta)=-\sum_{n=1}^{N}\ln\frac{m(1-\zeta)\cdot\mathbbm{1}(y_{n}\in\mathcal{M}_{1})+\zeta\cdot\mathbbm{1}(y_{n}\in\mathcal{M}_{2})}{k_{1}^{(n)}\cdot m(1-\zeta)+k_{2}^{(n)}\cdot\zeta}.$
Taking the derivative,
$f^{\prime}_{N,m}(\zeta)$ $=\sum_{n:y_{n}\in\mathcal{M}_{1}}\frac{1}{1-\zeta}-\sum_{n:y_{n}\in\mathcal{M}_{2}}\frac{1}{\zeta}-\sum_{n=1}^{N}\frac{mk_{1}^{(n)}-k_{2}^{(n)}}{(1-\zeta)mk_{1}^{(n)}+k_{2}^{(n)}\zeta}$
$=\sum_{n:y_{n}\in\mathcal{M}_{1},n\in\mathcal{I}}\frac{1}{1-\zeta}-\sum_{n\in\mathcal{I}:y_{n}\in\mathcal{M}_{2}}\frac{1}{\zeta}-\sum_{n\in\mathcal{I}}\frac{mk_{1}^{(n)}-k_{2}^{(n)}}{(1-\zeta)mk_{1}^{(n)}+k_{2}^{(n)}\zeta}$
$=\sum_{n\in\mathcal{I}:y^{(n)}\in\mathcal{M}_{1}}\left(\frac{1}{1-\zeta}-\frac{mk_{1}^{(n)}-k_{2}^{(n)}}{(1-\zeta)mk_{1}^{(n)}+k_{2}^{(n)}\zeta}\right)-\sum_{n\in\mathcal{I}:y^{(n)}\in\mathcal{M}_{2}}\left(\frac{1}{\zeta}+\frac{mk_{1}^{(n)}-k_{2}^{(n)}}{(1-\zeta)mk_{1}^{(n)}+k_{2}^{(n)}\zeta}\right).$
Let $\mathcal{I}_{1}=\{n\in\mathcal{I}\,|\,y^{(n)}\in\mathcal{M}_{1}\}$ and $\mathcal{I}_{2}=\{n\in\mathcal{I}|y^{(n)}\in\mathcal{M}_{2}\}$. From the first condition in the lemma statement, $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are non-empty. Note that
$\frac{mk_{1}^{(n)}-k_{2}^{(n)}}{(1-\zeta)mk_{1}^{(n)}+k_{2}^{(n)}\zeta}=\frac{1}{1-\zeta}-\frac{k_{2}^{(n)}}{(1-\zeta)^{2}mk_{1}^{(n)}+\zeta(1-\zeta)k_{2}^{(n)}},$
we can further simplify $f^{\prime}_{N,m}(\zeta)$ to
$f^{\prime}_{N,m}(\zeta)$ $=\sum_{n\in\mathcal{I}}\frac{k_{2}^{(n)}}{(1-\zeta)^{2}mk_{1}^{(n)}+\zeta(1-\zeta)k_{2}^{(n)}}-\sum_{n\in\mathcal{I}:y^{(n)}\in\mathcal{M}_{2}}\left(\frac{1}{\zeta}+\frac{1}{1-\zeta}\right).$
The derivative of the KL regularization $g(\zeta)=(1-\zeta)\ln\frac{1-\zeta}{p_{*}(1)}+\zeta\ln\frac{\zeta}{p_{*}(2)}$ is
$g^{\prime}(\zeta)=\ln\frac{\zeta}{1-\zeta}+\ln\frac{p_{*}(1)}{p_{*}(2)}<\frac{1}{1-\zeta}+\frac{1}{2\beta}|\{n\in\mathcal{I}\,|\,y^{(n)}\in\mathcal{M}_{2}\}|.$
Combining the above, we have
$f^{\prime}_{N,m}(\zeta)+\beta g^{\prime}(\zeta)$ $<\sum_{n\in\mathcal{I}}\frac{k_{2}^{(n)}}{(1-\zeta)^{2}mk_{1}^{(n)}+\zeta(1-\zeta)k_{2}^{(n)}}-\sum_{n\in\mathcal{I}:y_{n}\in\mathcal{M}_{2}}\frac{1}{1-\zeta}-\frac{1}{2}\sum_{n\in\mathcal{I}:y^{(n)}\in\mathcal{M}_{2}}\frac{1}{\zeta}+\frac{\beta}{1-\zeta}.$
@@
The first term is always positive and converges to 0 as $m\to\infty$ for a fixed $\zeta\in(0,1)$. Thus, there exists an $M>0$ such that for all $m>M$, $\sum_{n\in\mathcal{I}}\frac{k_{2}^{(n)}}{(1-\zeta)^{2}mk_{1}^{(n)}+\zeta(1-\zeta)k_{2}^{(n)}}<\frac{1}{2}\sum_{n\in\mathcal{I}:y^{(n)}\in\mathcal{M}_{2}}\frac{1}{\zeta}$. This implies that
$f^{\prime}_{N,m}(\zeta)+\beta g^{\prime}(\zeta)<\frac{\beta}{1-\zeta}-\sum_{n\in\mathcal{I}:y^{(n)}\in\mathcal{M}_{2}}\frac{1}{1-\zeta}<-\frac{1}{1-\zeta},$
where the last inequality follows from |$|\{n\in\mathcal{I}\,|\,y^{(n)}\in\mathcal{M}_{2}\}|>\max\{1+\beta,2\beta\ln\frac{p_{*}(1)}{p_{*}(2)}\}$. Thus, the overall loss function is always strictly decreasing at $\zeta\in(0,1)$ for large enough $m$. As a result, letting $\zeta_{m}^{*}=\operatorname*{arg\,min}f_{N,m}(\zeta)+g(\zeta)$, then $\lim_{m\to\infty}\zeta^{*}_{m}=1$.
See 6
@@
By Lemma 8, there exists a constant $\gamma\in(0,1)$ such that for all $N>0$, $\mathbb{P}(|\mathcal{I}|\leq N(1-\gamma))\leq e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}$. Let $q_{2}=\frac{1}{|\mathcal{I}|}\sum_{n\in\mathcal{I}}\mathbbm{1}(y_{n}\in\mathcal{M}_{2})$, and $\epsilon(\beta)=\max\{\beta+1,2\beta\ln\frac{p_{*}(1)}{p_{*}(2)}\}$, then $\mathbb{E}[q_{2}]=p_{*}(2)$. By Hoeffding’s inequality,
$\mathbb{P}\left(q_{2}-p_{*}(2)\leq\frac{1}{|\mathcal{I}|}\epsilon(\beta)\mid\mathcal{I}\right)<e^{-2|\mathcal{I}|\epsilon^{2}(\beta)}.$
Following the same argument using the law of total probability as in the proof for Lemma 9, we have
($\mathbb{P}\left(\left\{|\mathcal{I}|>N(1-\gamma)\right\}\cap\left\{\sum_{n\in\mathcal{I}}\mathbbm{1}(y_{n}\in\mathcal{M}_{2})>|\mathcal{I}|p_{*}(2)+\epsilon(\beta)\right\}\right)\geq 1-e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}-e^{-2N(1-\gamma)\epsilon^{2}(\beta)}.$
We let $\tilde{\delta}>2e^{-\frac{N(1-\gamma)\gamma^{2}\epsilon^{2}(\beta)}{8}}>e^{-\frac{N(1-\gamma)\gamma^{2}}{8}}+e^{-2N(1-\gamma)\epsilon^{2}(\beta)}$, which in turn gives $N>\frac{8\ln(2/\tilde{\delta})}{(1-\gamma)\gamma^{2}\epsilon^{2}(\beta)}$. Thus, for all $\delta>0$, for all $N>\frac{8\ln(2/\delta)}{(1-\gamma)\gamma^{2}\epsilon^{2}(\beta)}$, with probability $1-\delta$, we have
$|\mathcal{I}|>N(1-\gamma)\quad\text{and}\quad|\{n\in\mathcal{I}|y^{(n)}\in\mathcal{M}_{2}\}|>\max\left\{1+\beta,2\beta\ln\frac{p_{*}(1)}{p_{*}(2)}\right\}.$
We conclude the proof by applying Lemma 12.

### 21 Dec 2023 v2

@@ A.2 DPO
$P_{\pi_{r}}(x)=\frac{P_{\overline{\pi}}(x)e^{r(x)|\mathcal{D}|/\beta}}{\sum_{x^{\prime}\in\mathcal{M}}P_{\overline{\pi}}(x^{\prime})e^{r(x^{\prime})|\mathcal{D}|/\beta}}.$ (22)
@@
$-\sum_{(\mathcal{Y},y)\in\mathcal{D}}\left[\ln\left(\frac{e^{r(y)}}{\sum_{y^{\prime}\in\mathcal{Y}}e^{r(y^{\prime})}}\right)\right]=-\sum_{(\mathcal{Y},y)\in\mathcal{D}}\left[\ln\frac{\left(P_{\pi_{r}}(y)/P_{\overline{\pi}}(y)\right)^{\beta/|\mathcal{D}|}}{\sum_{y^{\prime}\in\mathcal{Y}}\left(P_{\pi_{r}}(y^{\prime})/P_{\overline{\pi}}(y^{\prime})\right)^{\beta/|\mathcal{D}|}}\right].$ (23)
Thus the value $\mathcal{L}_{\mathrm{reward}}(r)$ equals the value $\mathcal{L}_{\mathrm{DPO}}(\pi_{r})$. Note also that $r\mapsto\pi_{r}$ is surjective. Suppose $\widehat{r}$ is the minimizer for $\mathcal{L}_{\mathrm{reward}}$. Then $\pi_{\widehat{r}}$ is the optimal policy obtained by minimizing $\mathcal{L}_{\mathrm{policy}}(\cdot|\widehat{r})$. If $\pi_{\widehat{r}}$ is not optimal for $\mathcal{L}_{\mathrm{DPO}}$, then there exists another policy $\pi^{\prime}$ that obtains a strictly lower loss $\mathcal{L}_{\mathrm{DPO}}(\pi^{\prime})$. Since $r\mapsto\pi_{r}$ is surjective, there exists a reward function $r^{\prime}$ such that $\pi^{\prime}=\pi_{r^{\prime}}$. For example, we can take $r^{\prime}(x)=\frac{\beta}{|\mathcal{D}|}\ln\frac{P_{\pi^{\prime}}(x)}{P_{\overline{\pi}}(x)}$, which achieves a lower reward loss $\mathcal{L}_{\mathrm{reward}}(r^{\prime})=\mathcal{L}_{\mathrm{DPO}}(\pi_{r^{\prime}})<\mathcal{L}_{\mathrm{DPO}}(\pi_{\widehat{r}})=\mathcal{L}_{\mathrm{reward}}(\widehat{r})$, a contradiction. Similarly, if $\pi^{*}$ is optimal for $\mathcal{L}_{\mathrm{DPO}}$, the corresponding reward $r^{*}(x)=\frac{\beta}{|\mathcal{D}|}\ln\frac{P_{\pi^{*}}(x)}{P_{\overline{\pi}}(x)}$ must be optimal for $\mathcal{L}_{\mathrm{reward}}$. The corresponding optimal policy $\pi_{r^{*}}$ for $\mathcal{L}_{\mathrm{policy}}(\cdot|r^{*})$ then satisfies
$P_{\pi_{r^{*}}}(x)\propto P_{\overline{\pi}}(x)e^{\frac{\beta}{|\mathcal{D}|}\ln\frac{P_{\pi^{*}}(x)}{P_{\overline{\pi}}(x)}|\mathcal{D}|/\beta}=P_{\pi^{*}}(x),$
as desired. ∎
@@ A.3 Inclusive Learning
In this subsection, we prove Proposition 6. We start by defining two key sets that describe messages in $\mathcal{D}$. Recall that in the proof for Proposition 4, we defined
=$\mathcal{I}=\big{\{}(\mathcal{Y},y)\in\mathcal{D}\,|\,\mathcal{Y}\cap\mathcal{M}_{1}\neq\emptyset,\,\mathcal{Y}\cap\mathcal{M}_{2}\neq\emptyset\big{\}}.$
To facilitate the exposition, we also define
$\mathcal{I}_{\tau}=\left\{(\mathcal{Y},y)\in\mathcal{I}\mid y\in\mathcal{M}_{\tau}\right\}\quad\text{for }\tau=1,2$
to be the set of data in $\mathcal{I}$ where a message of category $\tau$ is chosen. We first prove the following lemma that characterizes the optimal solution for a class of convex loss functions.
@@
Suppose that for all $m>0$, the loss $\mathcal{L}(\cdot;m):(0,1)\to\Re$ is convex in $x$, $\mathcal{L}(x_{m};m)=\min_{x}\mathcal{L}(x;m)$, and there exists $M>0$ such that for all $m>M$,
$\frac{\partial\mathcal{L}(x;m)}{\partial x}<0.$
Then $\lim_{m\to\infty}x_{m}=1$.
Proof.
Since $\mathcal{L}(x;m)$ is convex on $(0,1)$, its local minima is also a global minima. For all $m>M$, $\mathcal{L}$ is strictly decreasing for $x\in(0,1)$, so $\lim_{m\to\infty}x_{m}=1$. ∎
Then, we define the high probability event
$\mathcal{E}(\mathcal{D})=\left\{|\mathcal{I}_{1}|>1\right\}\cap\left\{|\mathcal{I}_{2}|>\max\{1+\beta,2\beta\ln\frac{p_{*}(2)}{p_{*}(1)}\}\right\}.$
Lemma 14.
Under Assumption 3, for all fixed $\mathcal{M}_{1}$ and $\mathcal{D}$ such that $\mathcal{E}(\mathcal{D})$ holds, if $\widehat{\theta}$ minimizes $\mathcal{L}_{\mathrm{IL}}(\widehat{\pi}_{\theta})$, then
@@
First, we notice that $\mathcal{L}_{\mathrm{IL}}(\widehat{\pi}_{\theta})$ is strongly convex for $\beta>0$ and $\mathcal{L}_{\mathrm{IL}}(\widehat{\pi}_{\theta})\to\infty$ as $\|\theta\|_{2}\to\infty$, thus a minimizer for $\mathcal{L}_{\mathrm{IL}}(\widehat{\pi}_{\theta})$ exists and is unique. Let $m=\frac{|\mathcal{M}_{2}|}{|\mathcal{M}_{1}|}$ and $q=P_{\widehat{\pi}_{\theta}}(\mathcal{M}_{2})$. The loss function can be written in terms of $m$ and $q$ as
@@
$\mathcal{L}(q,m)=\mathcal{L}_{\mathrm{IL}}(\widehat{\pi}_{\theta})=-\sum_{(\mathcal{Y},y)\in\mathcal{D}}\ln\frac{m(1-q)\cdot\mathbbm{1}_{\mathcal{M}_{1}}(y)+q\cdot\mathbbm{1}_{\mathcal{M}_{2}}(y)}{m(1-q)\cdot|\mathcal{Y}\cap\mathcal{M}_{1}|+q\cdot|\mathcal{Y}\cap\mathcal{M}_{2}|}+\beta\left[(1-q)\ln\frac{1-q}{p_{*}(1)}+q\ln\frac{q}{p_{*}(2)}\right].$ (24)
Taking the partial derivative with respect to $q$, under the event $\mathcal{E}(\mathcal{D})$,
$\frac{\partial\mathcal{L}(q,m)}{\partial q}$ $=\sum_{(\mathcal{Y},y)\in\mathcal{I}}\frac{|\mathcal{Y}\cap\mathcal{M}_{2}|}{m(1-q)^{2}|\mathcal{Y}\cap\mathcal{M}_{1}|+q(1-q)|\mathcal{Y}\cap\mathcal{M}_{2}|}-\sum_{(\mathcal{Y},y)\in\mathcal{I}_{2}}\left(\frac{1}{q}+\frac{1}{1-q}\right)+\beta\left(\ln\frac{q}{1-q}+\ln\frac{p_{*}(1)}{p_{*}(2)}\right)$
<$\stackrel{{\scriptstyle(a)}}{{<}}\sum_{(\mathcal{Y},y)\in\mathcal{I}}\frac{|\mathcal{Y}\cap\mathcal{M}_{2}|}{m(1-q)^{2}|\mathcal{Y}\cap\mathcal{M}_{1}|+q(1-q)|\mathcal{Y}\cap\mathcal{M}_{2}|}-\frac{|\mathcal{I}_{2}|}{q}-\frac{|\mathcal{I}_{2}|}{1-q}+\frac{\beta}{1-q}+\frac{|\mathcal{I}_{2}|}{2}$
<$\stackrel{{\scriptstyle(b)}}{{<}}\sum_{(\mathcal{Y},y)\in\mathcal{I}}\frac{|\mathcal{Y}\cap\mathcal{M}_{2}|}{m(1-q)^{2}|\mathcal{Y}\cap\mathcal{M}_{1}|+q(1-q)|\mathcal{Y}\cap\mathcal{M}_{2}|}-\frac{|\mathcal{I}_{2}|}{2q}-\frac{1}{1-q},$
@@
where $(a)$ follows from $|\mathcal{I}_{2}|>2\beta\ln\frac{p_{*}(2)}{p_{*}(1)}$ and $\ln\frac{q}{1-q}<\frac{1}{1-q}$, and $(b)$ follows from $|\mathcal{I}_{2}|>1+\beta$ and $q<1$. For all $10<q<$ 1, the first term is always positive and converges to 0 as $m\to\infty$. Thus, there exists an $M>0$ such that for all $m>M$, $\sum_{(\mathcal{Y},y)\in\mathcal{I}}\frac{|\mathcal{Y}\cap\mathcal{M}_{2}|}{m(1-q)^{2}|\mathcal{Y}\cap\mathcal{M}_{1}|+q(1-q)|\mathcal{Y}\cap\mathcal{M}_{2}|}<\frac{|\mathcal{I}_{2}|}{2q}$. This implies that
$\frac{\partial\mathcal{L}(q,m)}{\partial q}<-\frac{1}{1-q}<0.$
The proof follows from applying Lemma 13. ∎
Finally, we prove that under the dichotomy data assumption, the event $\mathcal{E}(\mathcal{D})$ holds with high probability.
Lemma 15.
Under Assumption 1, if $|\mathcal{Y}|\geq 2$, then for fixed $\mathcal{M}_{1}$, $\mathbb{P}(\mathcal{E}(\mathcal{D}))\to 1$ as $|\mathcal{M}_{2}|\to\infty$.
@@
By Lemma 10, there exists a constant $\gamma\in(0,1)$ such that for all $|\mathcal{D}|>0$, $\mathbb{P}(|\mathcal{I}|\leq|\mathcal{D}|(1-\gamma))\leq e^{-\frac{|\mathcal{D}|(1-\gamma)\gamma^{2}}{8}}$. Let $\epsilon(\beta)=\max\{\beta+1,2\beta\ln\frac{p_{*}(1)}{p_{*}(2)}\}$, then $\mathbb{E}[\mathbbm{1}(y\in\mathcal{M}_{2})\mid(\mathcal{Y},y)\in\mathcal{I}]=p_{*}(2)$. By Hoeffding’s inequality,
$\mathbb{P}\left(\sum_{(\mathcal{Y},y)\in\mathcal{I}}\mathbbm{1}(y\in\mathcal{M}_{2})-|\mathcal{I}|\cdot p_{*}(2)\leq\epsilon(\beta)-|\mathcal{I}|\cdot p_{*}(2)\mid\mathcal{I}\right)$ $\leq e^{-2\frac{(\epsilon(\beta)-|\mathcal{I}|p_{*}(2))^{2}}{|\mathcal{I}|}}<e^{-2|\mathcal{I}|p_{*}^{2}(2)+4\epsilon(\beta)p_{*}(2)}$
$\mathbb{P}\left(\sum_{(\mathcal{Y},y)\in\mathcal{I}}\mathbbm{1}(y\in\mathcal{M}_{1})-|\mathcal{I}|\cdot p_{*}(1)\leq 1-|\mathcal{I}|\cdot p_{*}(1)\mid\mathcal{I}\right)$ $\leq e^{-2\frac{(1-|\mathcal{I}|p_{*}(2))^{2}}{|\mathcal{I}|}}<e^{-2|\mathcal{I}|p_{*}^{2}(2)+4p_{*}(2)}.$
Following a similar argument as that in the proof for Lemma 11, we have
$\mathbb{P}(\mathcal{E}(\mathcal{D}))$ ($\geq\mathbb{P}\left(\left\{|\mathcal{I}|>|\mathcal{D}|(1-\gamma)\right\}\cap\left\{|\mathcal{I}_{1}|>1\right\}\cap\left\{|\mathcal{I}_{2}|>\epsilon(\beta)\right\}\right)$
$\geq 1-e^{-\frac{|\mathcal{D}|(1-\gamma)\gamma^{2}}{8}}-e^{-2|\mathcal{D}|(1-\gamma)p_{*}^{2}(2)+4\epsilon(\beta)p_{*}(2)}-e^{-2|\mathcal{D}|(1-\gamma)p_{*}^{2}(2)+4p_{*}(2)}.$
The proof follows by taking $|\mathcal{M}_{2}|\to\infty$. ∎
Finally, we prove the following failure case for inclusive learning. See 6
Proof.
The proof follows from applying Lemmas 14 and 15, and noticing that $\mathbb{P}(\lim_{|\mathcal{M}_{2}|\to\infty}P_{\widehat{\pi}_{\widehat{\theta}}}(\mathcal{M}_{1})=0)\geq\mathbb{P}(\mathcal{E}(\mathcal{D}))$. ∎